I hope u can help me with this .
two angles are supplementary. one angle is 9degrees more than twice the other angle. Find angles.
thank u.


Answer by dbashant(7)

Angle x and angle y are supplementary, so x + y = 180.
angle x is 9 more than twice angle y, so x = 2y + 9.
substituing 2y + 9 for x in the first equation, we get (2y + 9) + y = 180.
Solve for y - 2y + 9 + y = 180 add the y's together
3y + 9 = 180 subtract 9 from each side
3y = 171 divide both sides by 3
y = 57
So, if x + y = 180 then x + 57 = 180 and x = 123
So x = 123. And y = 57.
To check - Is 123 9 more than twice 57? (2 X 57) + 9 = 114 + 9 = 123 CHECK!